# Evolutionary Generative Optimization: Towards Fully Data-Driven Evolutionary Optimization via Generative Learning

### v1

@@ IV Experiments
This section presents a comprehensive evaluation of EvoGO across diverse complex optimization tasks, including low- and high-dimensional numerical benchmarks, classical control problems, and complex high-dimensional robot control tasks. The experiments are designed to assess the effectiveness, efficiency, and scalability of our proposed data-driven evolutionary framework. Specifically, we aim to address the following research questions:
RQ1: How does EvoGO perform in terms of solution quality and convergence speed compared to representative baselines under standard and large population settings?
@@
By systematically exploring these questions, we validate the proposed framework through three groups of experiments: (1) performance and efficiency comparisons under standard and large-scale population regimes; (2) ablation studies to analyze the impact of different components and hyperparameters; and (3) visualizations to reveal the behavioral characteristics of the learned optimization process. These studies collectively demonstrate the scalability, adaptability, and data-efficiency of EvoGO in solving complex optimization problems without reliance on heuristic operations.
@@ IV-A Experimental Setup
Our empirical evaluation covers three categories of optimization challenges: (i) low- and high-dimensional numerical functions, (ii) low- and high-dimensional classical control tasks, and (iii) high-dimensional robotic control tasks. Specifically, the numerical benchmarks include functions from the CEC suite proposed in [90]. The classical control tasks “Pushing” and “Rover” are taken from [91], while “Landing”, “Walker”, and “Ant” are adapted from the Gym library [92]. Detailed descriptions of all benchmark tasks are provided in Supplementary Document C. The maximum number of fitness evaluations (FEs) is set to 1,000 for low-dimensional tasks and 10,000 for high-dimensional tasks.
EvoGO is compared against a comprehensive set of baseline methods, including: (i) Bayesian optimization algorithms with sequential or batch querying, such as Hyperband [93], TuRBO [54], and BOGP (LogEI) [48]; (ii) Tree-structured Parzen Estimator (TPE) [94], a widely used BO method in AutoML; (iii) evolutionary strategies and swarm intelligence algorithms, represented by CMA-ES [27] and PSO [26]; (iv) GPEME [65], a representative surrogate-assisted evolutionary algorithm; (v) Nelder–Mead [95], a classical heuristic method; and (vi) Random Search as a baseline.
All algorithms are evaluated over 20 independent runs with different random seeds on each benchmark. Further implementation details, including optimization settings and training configurations, are provided in Supplementary Document C, while ablation studies on key hyperparameters can be found in Supplementary Document G.
(a) Rosenbrock (5D)
@@
Figure 3: The convergence curves of EvoGO and the compared baselines in standard settings. The $y$-axes for numerical functions are in logarithmic scale, while those for other tasks are in linear scale since their true optimal is unknown. One-sigma error bars are adopted since two-sigma ones result in excessively large error bars for some algorithms. Please refer to Supplementary Document F for the complete figure.
(a) Ackley (200D)
(b) Rosenbrock (200D)
(c) Levy (200D)
(d) Walker (198D)
(e) Ant (404D)
(f) Plot Legend
Figure 4: The convergence curves of EvoGO and the compared baselines (TuRBO, CMA-ES, PSO, and Random Search) in a large-population setting (1000 FEs per generation). All configurations are consistent with those in Fig. 3, except that sequential acquisition algorithms are removed since due to unadjustable population sizes. Please refer to Supplementary Document F for complete figure.
@@ IV-B Convergence Performance across Benchmarks
To address our first research question (RQ1), we evaluate the performance of EvoGO against representative baselines. This set of experiments is designed to assess the final solution quality across a wide range of tasks. By testing on 21 benchmarks spanning numerical functions, classical control, and robot control, with dimensions from 5D to 404D, we aim to demonstrate the method’s broad applicability and effectiveness under both standard and large-population settings.
First, a subset of representative results under standard settings is shown in Fig. 3, while the full results are provided in Supplementary Document F. Under standard population settings, EvoGO consistently outperforms the baselines across most domains, including both low- and high-dimensional numerical functions (Figs. 3(a) to 3(d)), classical control tasks (Figs. 3(e) to 3(f)), and high-dimensional robot control problems (Figs. 3(g) to 3(h)). In total, EvoGO achieves the best performance on 19 out of 21 tasks, 10 of which are statistically significant, and matches the top-performing method on the remaining 9.
To assess scalability under unified conditions, we evaluate several high-dimensional benchmarks using a fixed population size of 1000 across all population-based baselines. As shown in Fig. 4, EvoGO consistently achieves the best performance, demonstrating its ability to efficiently exploit large population diversity in high-dimensional optimization tasks.
We further evaluate EvoGO on the Hopper environment provided by Brax [96], a high-dimensional robot control benchmark that enables massively parallel simulation through GPU acceleration. As shown in Fig. 5, EvoGO achieves the highest final reward among all compared methods, including TPE, CMA-ES, and PPO [97], with respective population sizes of 1000, 1, 100, and 64. Although PPO is an online reinforcement learning algorithm that continuously updates its policy during interaction with the environment, EvoGO achieves comparable or even superior performance. Moreover, EvoGO demonstrates a significant advantage in optimization efficiency: within a time budget of just 500 seconds, it consistently outperforms all baselines—including those running for over an hour. These results highlight the ability of our method to rapidly produce high-quality solutions in complex, hardware-accelerated continuous control tasks.
@@ IV-C Convergence Speed across Benchmarks
First, as summarized in Table I, EvoGO converges within 10 generations across all large-scale benchmark tasks. In contrast, baseline methods often require tens to thousands of generations to achieve comparable performance. This property is crucial for scenarios with tight runtime budgets but abundant parallel resources, where large populations can be evaluated concurrently.
TABLE I: The number of generations to complete each benchmark by each method. Note: sequential methods are not listed here since they use exactly the same number of generations as FEs.
@@
Numerical 10$\sim$100 83$\sim$526 25$\sim$50 956$\sim$9395 10
Landing 10 91 25 823$\sim$901 10
Pushing 10 91 25 862$\sim$933 10
Rover 100 625 25 3163$\sim$3216 10
Walker 100 526 50$\sim$51 1833$\sim$1839 10
Ant 100 455 24$\sim$25 1500$\sim$1503 10
Second, the runtime-constrained setting of Brax allows us to evaluate the wall-clock efficiency of each method under the same budget of 10,000 function evaluations and one hour of runtime. As illustrated in Fig. 5, EvoGO rapidly approaches a reward of 900 within the first 100 seconds, while other methods take significantly longer to reach similar performance—or fail to reach it at all. In particular, EvoGO achieves up to 134×134$\times$134 × speedup over CMA-ES and 50×50$\times$50 × over TPE in terms of reward per second. This demonstrates the strong convergence speed of our framework when leveraging large population sizes and parallel evaluations.
@@ IV-D Ablation Studies
To answer our second research question (RQ2), we conduct a comprehensive ablation study to dissect EvoGO and quantify the contribution of its key components. These experiments are designed to isolate the impact of elements like the surrogate model, infill criterion, and generative architecture on overall performance.
@@
Figure 7: The convergence curves of EvoGO and its ablations. “EvoGO” represents the original method; “Single Net” excludes the generative model $\bm{\gamma}^{\prime}$ and related loss terms; “Cycle GAN” adapts the Cycle GAN; “EvoGO-MLP” utilizes MLP for surrogate modeling; “EvoGO-LCB” employs the LCB infill criteria. Please refer to Supplementary Document G for the complete figure.
1) Upper Bound of the Generative Model
We begin by investigating the maximum potential of our framework to understand its upper performance bound. This experiment assesses the effectiveness of the overall architecture, independent of any surrogate model error. To achieve this, we replace the surrogate model’s predictions with ground-truth function evaluations in the optimization loss (“Real Eval”). This design serves two purposes: first, it reveals the high potential of our composite generative structure and optimization strategy, and second, it suggests that performance could be further enhanced with more accurate surrogate models. As presented in Fig. 6, the “Real Eval” method shows robust and consistent performance across various numerical functions, illustrating the effectiveness of EvoGO’s overall architecture. The great potential of the proposed composite generative model structure and the optimization loss is well demonstrated by the results, which can be considered as an upper bound of EvoGO.
2) Alternative Surrogate Models
We explore the framework’s modularity and sensitivity to the choice of surrogate model. This experiment is designed to understand how performance is influenced by different modeling assumptions. We replace the default Gaussian Process (GP) with a Multi-Layer Perceptron (MLP) [98] (“EvoGO-MLP”), a common alternative that lacks inherent uncertainty quantification. In both variants, the standard deviation and covariance functions are removed from the acquisition loss (Eq. 11). This allows us to assess the importance of the GP’s probabilistic nature and to demonstrate the framework’s adaptability. As shown in Fig. 7, the “EvoGO-MLP” struggled with some numerical functions, likely due to overfitting and lacking the variance estimation provided by GPs. However, it exhibited impressive results on complex scenarios like the Ant benchmark. These findings highlight EvoGO’s adaptability and its potential utility with different surrogate modeling methods.
We examine the modular structure of the framework and its performance under different surrogate modeling assumptions. We replace the default Gaussian Process (GP) with a Multi-Layer Perceptron (MLP) [98] (referred to as “EvoGO-MLP”), a common alternative that lacks inherent uncertainty quantification. In both variants, the standard deviation and covariance components are removed from the acquisition loss (Eq. 11), enabling an evaluation of the role of probabilistic modeling in GP-based surrogates and demonstrating the framework’s flexibility. As shown in Fig. 7, “EvoGO-MLP” underperforms on certain numerical benchmarks, likely due to overfitting and the absence of variance estimation available in GPs. However, it performs well on complex control tasks such as the Ant benchmark. These results demonstrate that EvoGO remains effective across different surrogate modeling choices while highlighting the value of uncertainty modeling in challenging scenarios.
3) Alternative Infill Criteria
We replace our tailored loss (Eq. 11) with one derived from the widely used Lower/Upper Confidence Bound (LCB/UCB) infill criteria (“EvoGO-LCB”) to validate the effectiveness of our custom optimization loss. By comparing the performance of variant against the original EvoGO, we can determine the specific benefits of criterion and test the framework’s flexibility with alternative infill criteria. Specifically, the optimization loss of the LCB/UCB infill criteria can be written as:
@@
where the derivation is detailed in Supplementary Document A. As shown in Fig. 7, the “EvoGO-LCB” variant generally performed worse than the original EvoGO. However, it still achieved competitive results on the robot control and some numerical benchmarks. This not only suggests that the original optimization loss of EvoGO (Eq. 11) is more robust across a wider range of tasks but also highlights the adaptability of the proposed EvoGO.
@@
4) Without Surrogate Model
In this section, we evaluate whether the surrogate-guided acquisition process is essential, or if the generative component alone can effectively guide the search. To this end, we test two variants: one that completely removes the surrogate model and its acquisition loss (“No Surrogate”), and another that replaces it with a standard adversarial loss from a Cycle GAN framework [75] (“Cycle GAN”). Specifically, the latter incorporates the data setup described in Algorithm 1, and replaces GP posterior and related acquisition loss (Eq. 11) with discriminative model and adversarial loss, respectively. Moreover, to train the “Cycle GAN” variant, we adopted the gradient inversion technique, which is not present in EvoGO. This allows us to evaluate the necessity of the surrogate model for balancing exploration and exploitation. As shown in Fig. 7, these models generally showed diminished effectiveness, likely due to increased susceptibility to mode collapse. In contrast, EvoGO adaptively balances exploration and exploitation, thus avoiding such pitfalls.
5) Unpaired Generative Modeling
Finally, we investigate the importance of the paired generative architecture, a core structural choice in EvoGO. To do this, we test a “Single Net” variant with only one generative model, removing the losses that enforce consistency between the pair (Eq. 5). Specifically, it eliminates the reconstruction loss $\mathcal{L}_{\text{rec}}$ (Eq. 5) and half of the similarity loss $\mathcal{L}_{\text{sim}}$ (Eq. 6) in Eq. 7. Furthermore, the standard deviation loss $\mathcal{L}_{\text{std}}$ in Eq. 11 is also removed in the “Single Net” variant. This experiment is crucial for validating our hypothesis that the paired design is key to preventing mode collapse and ensuring robust performance. The results shown in Fig. 7 confirms that this variant underperformed significantly across most benchmarks, underscoring the critical role of paired modeling in maintaining population diversity and enhancing overall system performance.
6) Extended Ablations
Additional ablation studies are detailed in Supplementary Document G. These studies examined variations in the neural network architectures and hyperparameter configurations, further elucidating the robustness and adaptability of EvoGO across various settings. Specifically, we explored the impact of different generative model layers and the width of them, noticing that the generative model’s architecture has little effect on EvoGO’s performance, as shown in Fig. S4. Moreover, we also investigated the impact of the reconstruction loss scale $\lambda_{1}$, the data split factor $\eta$, and the sliding window factor $\epsilon$, all of which were found to have small to little impact on the performance of EvoGO, as shown in Figs. S5 to S7.

### v2

@@ IV Experiments
This section presents a comprehensive evaluation of EvoGO across diverse complex optimization tasks, including low- and high-dimensional numerical benchmarks, classical control problems, and complex high-dimensional robot control tasks. The experiments are rigorously designed to assess the effectiveness, efficiency, and scalability of our proposed data-driven evolutionary framework. Specifically, we aim to address the following research questions:
RQ1: How does EvoGO perform in terms of solution quality and convergence speed compared to representative baselines?
@@
To systematically address these research questions, we validate the proposed framework through three cohesive groups of experiments: (1) performance and efficiency comparisons across complex black-box optimization problems with varying dimensionalities; (2) ablation studies to analyze the impact of different components and hyperparameters; and (3) visualizations to reveal the behavioral characteristics of the learned optimization process. These studies collectively demonstrate the scalability, adaptability, and data-efficiency of EvoGO in solving complex optimization problems without reliance on heuristic operations.
@@ IV-A Experimental Setup
Our empirical evaluation covers three categories of optimization challenges: (i) low- and high-dimensional numerical functions, (ii) low- and high-dimensional classical control tasks, and (iii) robotic control tasks. The classical control tasks Pushing and Rover are taken from [95], while Landing, Walker, and Ant are adapted from the Gym library [96]. The maximum number of fitness evaluations (FEs) is set to 1,000 for low-dimensional tasks and 10,000 for high-dimensional tasks.
EvoGO is compared against a comprehensive set of baseline methods, including: (i) Bayesian optimization algorithms, covering both sequential and batch variants, such as Hyperband [97], TuRBO [56], SCBO [98], BOGP (LogEI) [50], and the Tree structured Parzen Estimator (TPE) [99], which is widely used in AutoML; (ii) evolutionary strategies, represented by CMA-ES [11]; (iii) Nelder Mead [100], a classical heuristic method; (iv) SADE-SS [101] and SA-LSEO-LE [102], two state of the art surrogate assisted evolutionary algorithms.
All algorithms are evaluated over 20 independent runs with different random seeds on each benchmark. Additional details regarding the experimental environment, parameter configurations, and parallel execution settings are provided in Supplementary Document C, while ablation studies on key hyperparameters are included in Supplementary Document F.
(a) Rosenbrock (5)
@@
Figure 3: The convergence curves of EvoGO and the compared baselines in standard settings. The $y$-axes for numerical functions are in logarithmic scale, while those for other tasks are in linear scale since their true optimal is unknown. One-sigma error bars are adopted since two-sigma ones result in excessively large error bars for some algorithms. Please refer to Supplementary Document E for the complete figure.
@@ IV-B Convergence Performance across Benchmarks
To address our first research question (RQ1), we evaluate the performance of EvoGO against a comprehensive set of representative baselines. This critical set of experiments is designed to rigorously assess the final solution quality and convergence speed across diverse tasks. Specifically, by testing on 25 benchmarks spanning numerical functions, classical control, and robot control, with dimensions ranging from 10D to 1000D, we aim to demonstrate the broad applicability and effectiveness of the method across various dimensionalities.
TABLE S.I: Statistical Results (Mean and Standard Deviation) of the Tested Algorithms on 1000-Dimensional Benchmark Problems. Each Algorithm Was Executed Ten Times, and the Best Results Are Highlighted in Bold.
Problem CMA-ES Nelder-Mead TuRBO SA_LSEO_LE EvoGO
Ackley 2.1192e+01 (2.66e-02) 2.1228e+01 (8.64e-03) 2.1236e+01 (2.80e-02) 2.0169e+01 (1.37e-01) 1.3652e+01 (1.10e-01)
Rosenbrock 1.4956e+09 (1.55e+08) 4.4097e+09 (1.73e+08) 5.9143e+08 (1.93e+07) 3.6834e+07 (1.32e+07) 5.8708e+06 (3.50e+05)
Rastrigin 9.2663e+05 (2.17e+04) 1.1251e+06 (1.73e+04) 4.2006e+05 (8.88e+03) 7.6479e+04 (7.78e+03) 5.4188e+04 (5.74e+02)
Levy 5.4255e+05 (3.66e+04) 5.8105e+05 (1.34e+04) 2.2472e+05 (7.57e+03) 1.6155e+05 (1.35e+04) 2.4058e+04 (8.77e+02)
The aggregated results, a representative subset of which is shown in Fig. 3 (with full details in Supplementary Document E), demonstrate that EvoGO consistently outperforms the baselines across most domains. A total of 25 benchmark tasks are evaluated, including four 1000-dimensional numerical problems listed in Table S.I. Due to the long runtime of high-dimensional cases, one representative comparison is provided for each algorithm. Overall, on low-dimensional problems, our method outperforms all baselines except the state-of-the-art SAEA algorithm SADE-SS, which can better exploit limited data in small-sample settings. However, for high-dimensional problems (e.g., 1000D), SADE-SS requires over one day of computation for a single run, making it impractical for large-scale applications. As shown in Table S.I, our method surpasses SA-LSEO-LE (state-of-the-art high-dimensional SAEA variants), indicating its superior capability to learn underlying regularities and distributions in complex black-box optimization tasks.
We further evaluate EvoGO on the Hopper environment provided by Brax [103], a high-dimensional robotic control benchmark that supports massively parallel simulation via GPU acceleration. As shown in Fig. 4, EvoGO achieves the highest final reward among all compared methods, including TPE, CMA-ES, and PPO [104], with respective population sizes of 1000, 1, 100, and 64. Notably, EvoGO attains comparable or even higher performance than PPO, despite the latter being an online reinforcement learning algorithm that continuously updates its policy through environment interaction. Furthermore, EvoGO exhibits substantial optimization efficiency: within a time budget of only 500 seconds, it surpasses all baselines, including those that continue training for over an hour. These results underscore the capability of EvoGO to deliver rapid and high-quality solutions in complex, GPU accelerated continuous control environments.
@@ IV-C Convergence Speed across Benchmarks
First, as summarized in Table S.II, EvoGO converges within 10 generations across all large-scale benchmark tasks. In contrast, baseline methods often require tens to thousands of generations to achieve comparable performance. This property is crucial for scenarios with tight runtime budgets but abundant parallel resources, where large populations can be evaluated concurrently.
TABLE S.II: The number of generations to complete each benchmark by each method. Note: sequential methods are not listed here since they use exactly the same number of generations as FEs.
@@
[…]
Ant $4\sim25 1500\sim$1503 10
Second, the runtime constrained setting of Brax allows us to evaluate the wall clock efficiency of each method under the same budget of 10,000 function evaluations and one hour of runtime. As shown in Fig. 4, EvoGO rapidly attains a reward of 900 within the first 100 seconds, while other methods require substantially more time to reach comparable performance, and some never achieve it. In particular, EvoGO delivers up to $134\times$ higher reward per second than CMA ES and $50\times$ higher than TPE, reflecting its superior optimization throughput. These findings highlight the robust convergence rate of EvoGO when utilizing large population sizes and parallel evaluations.
@@ IV-D Ablation Studies
To answer our second research question (RQ2), we conduct a comprehensive ablation study to dissect EvoGO and quantify the contribution of its key components. These experiments are designed to isolate the impact of elements such as the surrogate model, infill criteria, and generative architecture on overall performance.
@@
Figure 6: The convergence curves of EvoGO and its ablations. “EvoGO” represents the original method; “Single Net” excludes the generative model $\bm{\gamma}$ and related loss terms; “Cycle GAN” adapts the Cycle GAN; “EvoGO-MLP” utilizes MLP for surrogate modeling; “EvoGO-LCB” employs the LCB infill criteria. Please refer to Supplementary Document F for the complete figure.
1) Under Perfect Surrogate Model
To examine the performance limit of EvoGO, we analyze its behavior when the surrogate model is assumed to be perfectly accurate. In this setting, the predicted objective values in the optimization loss are replaced with the ground-truth function evaluations, forming a variant referred to as Real Eval. This experiment isolates the effectiveness of the overall generative architecture and optimization strategy from any surrogate-induced error, thereby assessing the theoretical upper bound of the framework. As shown in Fig. 5, the Real Eval configuration achieves robust and consistent results across different numerical benchmarks, demonstrating the full potential of the proposed composite generative structure and loss formulation. These results indicate that EvoGO can achieve even higher performance when the surrogate model approaches perfect accuracy.
2) Under Failed Surrogate Model
To simulate the scenario where the surrogate model completely fails to provide reliable guidance, we evaluated two variants of EvoGO that remove or replace the surrogate component. The first variant, denoted as No Surrogate, eliminates the surrogate model and its corresponding optimization loss, relying solely on the generative mechanism for search guidance. The second variant, referred to as CycleGAN, replaces the surrogate-based optimization loss with a standard adversarial loss from the CycleGAN framework [80]. These two configurations effectively emulate conditions in which the surrogate model is either absent or incapable of meaningful prediction. As shown in Fig. 6, both variants exhibited noticeable performance degradation and increased instability, indicating that the surrogate model plays a crucial role in maintaining effective exploration-exploitation balance. Nevertheless, the generative component still preserved certain diversity and structural stability, largely due to the reconstruction constraint that prevents severe mode collapse. These observations confirm that while EvoGO can remain functional under surrogate failure, accurate surrogate feedback is vital for achieving its full performance potential.
3) Alternative Surrogate Model
To examine whether the proposed framework depends exclusively on GP surrogates, we conduct two complementary studies that evaluate both model substitution and infill criteria. These analyses reveal how EvoGO behaves when the GP-based probabilistic assumptions are replaced by alternative designs.
MLP as surrogate model. To evaluate the modularity of the surrogate component, we replaced the default GP with a Multi-Layer Perceptron (MLP) [105], forming the variant EvoGO-MLP. Since the MLP does not provide inherent uncertainty estimation, we removed the standard deviation and covariance terms from the optimization loss (Eq. 11). This configuration isolates the contribution of the probabilistic structure introduced by the GP. As shown in Fig. 6, the EvoGO-MLP variant showed slightly weaker performance on several numerical benchmarks, likely attributable to the absence of variance-guided exploration, while achieving strong results in complex scenarios such as the Ant benchmark. These results demonstrate that EvoGO remains effective even when the surrogate does not include explicit uncertainty modeling, highlighting the adaptability of the proposed framework.
infill criteria as replacement. We investigated the impact of the infill criteria by examining whether the proposed mechanism for balancing exploration and exploitation could be replaced by an alternative criterion. Specifically, we substituted the proposed optimization loss (Eq. 11) with the classical Lower/Upper Confidence Bound (LCB/UCB) criterion, thereby establishing the variant designated as EvoGO-LCB. The corresponding loss function is defined as
@@
The derivation for the LCB infill criteria is provided in Supplementary Document B. As illustrated in Fig. 6, the “EvoGO-LCB” variant achieved competitive performance but generally performed worse than the original formulation. This observation indicates that, while the LCB/UCB-based infill criteria maintains reasonable performance, the tailored optimization loss employed in EvoGO provides a more stable exploration mechanism across diverse problem settings.
@@
4) Paired Architecture Analysis
We further analyze the structural importance of the paired generative architecture, which is a fundamental component of EvoGO. To this end, we design a “Single Net” variant that removes one generator and all consistency-enforcing mechanisms between $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$. Specifically, this variant eliminates the reconstruction loss $\mathcal{L}_{\text{rec}}$ (Eq. 5) and one branch of the similarity loss $\mathcal{L}_{\text{sim}}$ (Eq. 6) in Eq. 7, while omitting the standard deviation regularization term $\mathcal{L}_{\text{std}}$ in Eq. 11. This ablation assesses whether a single-generator design can sustain the stability and diversity achieved through paired modeling. As presented in Fig. 6, the “Single Net” configuration leads to a noticeable degradation in convergence stability and diversity maintenance, particularly under complex multimodal settings. These findings demonstrate that the paired architecture, reinforced by the reconstruction constraint, is essential for ensuring balanced training dynamics, preventing mode collapse, and enhancing the robustness of the overall generative process.
5) Extended Ablations
Additional ablation studies are detailed in Supplementary Document F. These studies examined variations in the neural network architectures and hyperparameter configurations, further elucidating the robustness and adaptability of EvoGO across various settings. Specifically, we explored the impact of different generative model layers and their width. We observed that the generative model’s core architecture has little effect on EvoGO’s performance, as shown in Fig. S4. Moreover, we also investigated the impact of the reconstruction loss scale $\lambda_{1}$, the data split factor $\eta$, and the sliding window factor $\epsilon$. These components were all found to have small or negligible impact on the performance of EvoGO, as shown in Figs. S5 to S7.